Suppose that 5 years from now you will receive $10,000 at the end of every year for 5 years. What is the present value of this annuity if the opportunity cost rate is 5%?
Answer #1
The discounted value of the 5 year annuity at t4 = 10000*(1.05^5-1)/(0.05*1.05^5) = $       43,294.77
PV of the annuity at t0 = 43294.77/1.05^4 = $       35,618.71